# MAYA: Addressing Inconsistencies in Generative Password Guessing through a Unified Benchmark

### v2

@@ Abstract
Recent advances in generative models have led to their application in password guessing, with the aim of replicating the complexity, structure, and patterns of human-created passwords. Despite their potential, inconsistencies and inadequate evaluation methodologies in prior research have hindered meaningful comparisons and a comprehensive, unbiased understanding of their capabilities. This paper introduces MAYA, a unified, customizable, plug-and-play benchmarking framework designed to facilitate the systematic characterization and benchmarking of generative password-guessing models in the context of trawling attacks. Using MAYA, we conduct a comprehensive assessment of six state-of-the-art approaches, which we re-implemented and adapted to ensure standardization. Our evaluation spans eight real-world password datasets and covers an exhaustive set of advanced testing scenarios, totaling over $15,000$ compute hours. Our findings indicate that these models effectively capture different aspects of human password distribution and exhibit strong generalization capabilities. However, their effectiveness varies significantly with long and complex passwords. Through our evaluation, sequential models consistently outperform other generative architectures and traditional password-guessing tools, demonstrating unique capabilities in generating accurate and complex guesses. Moreover, the diverse password distributions learned by the models enable a multi-model attack that outperforms the best individual model. By releasing MAYA, we aim to foster further research, providing the community with a new tool to consistently and reliably benchmark generative password-guessing models. Our framework is publicly available at https://github.com/williamcorrias/MAYA-Password-Benchmarking.
@@ 1 Introduction
Beyond these limitations, traditional Machine Learning password guessing techniques, such as Probabilistic Context-Free Grammars (PCFGs) [31, 12, 63, 24] and Markov Models [38, 15, 22, 62], generate passwords based on statistical likelihood derived from real-world observed data. While these methods have demonstrated considerable effectiveness, they are fundamentally constrained by structural assumptions. Specifically, Markov Models, which rely on fixed-length n-grams, are inherently limited when modeling long-range dependencies and complex patterns. Likewise, PCFGs impose rigid template structures that constrain the diversity and flexibility of the generated password space.
@@
In recent years, research has increasingly focused on harnessing advancements in generative models to enhance password guessing techniques [29, 45, 42, 48, 40, 65, 61, 55]. Unlike traditional approaches, these models aim to learn and replicate the complexity, structure, and patterns of human-created passwords without relying on prior assumptions. Nevertheless, the current literature suffers from inconsistencies, inadequate evaluation practices, and a lack of rigorous, standardized characterization of generative models’ behaviors. Proposed approaches are often evaluated using inconsistent methodologies across studies and are tested in non-uniform, often simplistic settings, hindering meaningful comparisons and limiting a full understanding of their capabilities. As a result, critical aspects remain underexplored: how models behave across different settings, what types of passwords they tend to generate, how human-like their guesses are, whether they genuinely capture the underlying complexities of human password behavior, and where potential blind spots may lie. These limitations underscore a broader absence of systematic model characterization.
@@
To address these gaps, we introduce MAYA  [^1] The name ”MAYA ” is inspired by A. Schopenhauer’s philosophical concept of ”the veil of Maya”, representing an illusion that obscures true reality. Similarly, our framework aims to unveil the potential of password-guessing generative models, which has remained obscured until now. , a unified and customizable plug-and-play benchmarking framework. While MAYA is specifically designed to support fair comparison and in-depth characterization and evaluation of generative password-guessing models in trawling attack scenarios, it builds upon a general, standardized, and rigorous evaluation methodology that can be applied across a wide range of guessing attacks, regardless of the underlying guessing approach. Using MAYA, we systematically evaluate six state-of-the-art generative models, each re-implemented with standardized data preprocessing, dependencies, and configurations. We focus on trawling password attacks, as they represent the most general and widely applicable setting. Our experimental setup spans eight real-world password datasets for training and testing, and covers a wide range of evaluation scenarios, totaling over $15,000$ hours of computation.
@@
Our key findings include: (a) Increasing password length yields diminishing returns in reducing guessability; similarly, increasing the number of generated passwords offers reducing improvements in successful guesses, as models tend to exhaust easily guessable passwords early in the generation process. (b) Rule-based tools are competitive on smaller datasets and traditional machine learning models excel on challenging ones. However, the best-performing generative models, on average, outperform both. (c) Generative models do not always require large datasets to effectively model data distribution. While some architectures benefit from larger datasets, others exhibit minimal improvement or even a decrease in performance as the training data increases. (d) Models can successfully guess passwords even when trained on datasets from different communities and/or cultures compared to the test set, highlighting strong generalization capabilities and suggesting the existence of common structures underlying human-created passwords across disparate groups. (e) While all models demonstrate strong capabilities in guessing common and simple passwords and a decrease in performance for rarer passwords, sequential models are the only ones that remain effective in guessing longer passwords and more complex patterns. (f) Different architectures learn generation functions with different probability distributions over the codomain, enabling their combination into a multi-model attack that outperforms individual models. (g) Generative models effectively capture various aspects of human-created passwords, generating high-quality and diverse passwords while minimizing mode failures. However, some models struggle to accurately replicate the length distribution.
@@
  • We develop MAYA, a fully customizable, plug-and-play framework for evaluating generative password-guessing models in trawling attack scenarios. To support reproducibility and encourage further research, we publicly release our code and data at https://github.com/williamcorrias/MAYA-Password-Benchmarking.
  • Leveraging MAYA, we characterize and benchmark six state-of-the-art generative models across eight datasets and a diverse set of testing scenarios, addressing seven key research questions. Our evaluation spans over 15,000 compute hours.
@@ 2.1 Lack of Consistency
Each existing approach adopts its own evaluation methodology, as models are neither trained nor assessed on the same data and under identical settings. These methodological inconsistencies encompass various factors, including data preprocessing algorithms, file encoding, vocabulary, maximum password length, the number of generated passwords, and the size of the training and testing datasets. As a result, fair comparisons across studies are challenging. For instance, differing data preprocessing methods lead to distinct training and testing distributions, making direct comparison challenging. Likewise, evaluating models with different maximum password lengths introduces inherent bias, as shorter passwords are generally easier to guess. Analogous considerations apply to other settings as well. Such inconsistencies highlight the necessity for a standardized evaluation methodology to serve as a foundation for future research.
@@ 2.2 Lack of Rigorousness
Existing research often suffers from an insufficiently rigorous evaluation. Models are typically evaluated using overly simplistic metrics, such as the percentage of guessed passwords or the number of unique passwords generated, within restricted and similarly simplistic scenarios, often relying only on widely used datasets like RockYou or LinkedIn. Consequently, the current literature offers an incomplete evaluation that is biased toward these simplistic settings, underscoring the need for more robust methodology incorporating diverse datasets, varied scenarios, and more complex settings that reflect a broader spectrum of real-world contexts.
@@ 2.4 Why Trawling Attacks
Password attacks are typically categorized based on the adversary’s scope: targeted and trawling. Targeted attacks aim to compromise specific user accounts by leveraging personal identifiable information (PII), and have gained increasing attention in recent years due to their growing effectiveness [35, 61, 43, 60, 27, 64]. In contrast, trawling attacks seek to guess as many user passwords as possible within a dataset, without targeting any specific individual. This attack model has long been the focus of password security research, with numerous techniques developed over time, including rule-based systems, PCFGs, Markov models, traditional neural networks, and deep generative models. In this work, we focus exclusively on trawling attacks, as they represent the most general and widely applicable class of password-guessing attacks, providing a broad evaluation framework. Moreover, since targeted attacks can be viewed as trawling attacks conditioned on PII, our evaluation methodology can effectively support both lines of research.
@@ 3 MAYA
This section presents MAYA, our unified and plug-and-play benchmarking framework tailored for password-guessing generative models in the context of trawling attacks. The framework offers an intuitive environment for training and testing generative models with minimal setup, enabling the research community to move in a unified direction. It also includes a comprehensive set of experimental settings, providing thorough benchmarking and detailed characterization across multiple key metrics. MAYA features a highly modular and easily extendable architecture that enables the integration of new models and customized testing scenarios to support future research. While MAYA focuses on generative models, its underlying evaluation methodology (Section 3.1) provides a standardized testing environment that is applicable to all types of trawling password-guessing methods, ranging from traditional techniques to the latest approaches. MAYA currently implements six state-of-the-art password-guessing generative models (Section 3.2), eight real-world password datasets (Section 4), and a comprehensive set of testing scenarios aimed at answering seven key research questions (Section 3.3).
@@ 3.1.1 Standardized Data Preprocessing and Settings
We propose the following procedure to standardize data preprocessing: (1) open and read datasets using UTF-8 encoding while ignoring errors, (2) remove passwords that exceed the specified maximum length, contain non-ASCII characters, or include characters that are not in the provided vocabulary. (3) split the dataset following the standard $80\%$ training and $20\%$ testing ratios, (4) remove duplicates from the testing dataset, and (5) remove from the training dataset any overlap with the testing dataset. This approach ensures wide language compatibility through UTF-8 encoding, avoids double-counting by eliminating duplicates, and provides a fixed, consistent testing set across experiments. We note that existing works typically define the training set first, and then derive the testing set by removing any overlapping samples. This approach is undesirable, as varying the training set leads to changes in the testing set as well, thereby limiting the comparability across experiments. We further define our vocabulary to include all uppercase and lowercase letters, digits, and the following widely-accepted symbols: ~!@#\$%^&*(),.<>/?’"{}[]\-_=+;: ‘.
@@ 3.2 Models
We selected six state-of-the-art generative models for trawling password guessing, aiming to provide a comprehensive evaluation of different architectural approaches: (1) FLA [40], based on an LSTMs, (2) PassGAN [29], and (3) PLR-GAN [45], both GAN-based approaches, (4) PassFlow [42], a flow-based model, (5) VGPT2 [7], which combines a VAE with GPT2-derived encoder and decoder blocks, and (6) PassGPT [48], an autoregressive transformer model based on the GPT2 architecture. To ensure standardization, each model was adapted and re-implemented in strict accordance with the procedures described in the respective papers, minimizing discrepancies arising from implementation variations. We provide further implementation details in Appendix A. We validated the accuracy of our implementations by comparing our results with those reported in the original papers.
@@
FLA [40] (Fast, Lean, and Accurate), was the first approach to apply neural networks to the password-guessing task. It is the only approach we selected based on a recurrent neural network, specifically an LSTM, allowing us to examine how models designed for sequential data processing perform compared to other, more recent generative architectures.
@@
PLR-GAN [45] is an enhanced version of PassGAN and represents the current state-of-the-art for GAN-based methods. PLR-GAN further adopts a Dynamic Password Guessing (DPG) strategy, which allows the model to adapt its guesses based on the distribution of successfully guessed passwords, increasing the likelihood of generating relevant guesses.
@@
PassGPT [48] proposes a GPT-2-based language model for password generation. Similarly to FLA, PassGPT employs a sequential generation process. However, GPT-2 is built upon an attention mechanism, which allows it to capture long-range dependencies more effectively.
@@ 3.3 Research Questions
RQ2 - Are Generative Models Truly Better than Traditional Tools?
Despite recent advancements in generative architectures, there is limited evidence indicating whether generative models outperform traditional approaches, and few direct experimental comparisons have been conducted. We address this gap by conducting a comprehensive evaluation of generative models on 8 different datasets and comparing them to traditional tools such as John the Ripper (JtR) and Hashcat, offering a thorough, direct comparison between generative and rule-based methods.
@@
In real-world scenarios, attackers often have limited or partial access to leaks. Therefore, evaluating model performance across varying training data subset sizes is essential for gaining a clear understanding of the effectiveness of generative models. This RQ examines the ability of generative models to reconstruct the full target data distribution starting from varying portions of the source dataset and execute a successful attack. We explore this capability by training models on up to seven different data subsets on four distinct datasets.
@@
Simple passwords are easily guessed using traditional tools, whereas rare and complex passwords present a significantly greater challenge. The ability of generative models to arbitrarily sample from different areas of the password distribution offers the potential to guess even rare and complex passwords. We explore this RQ by dividing and categorizing the test dataset based on password frequency and assessing each model’s performance on the different subsets. Additionally, we analyze how well the models perform in guessing passwords of varying lengths and patterns.
@@
We explore the first part of this RQ using two metrics computed across all model pairs $(M_{1},M_{2})$: Jaccard Index and Mergeability Index. Jaccard Index measures the ratio between the intersection and union of the passwords generated by each model, providing a metric for the overall diversity of all passwords generated by $M_{1}$ and $M_{2}$. Additionally, we introduce the Mergeability Index, which quantifies the ratio between the marginal gain achieved by combining the outputs of $M_{1}$ and $M_{2}$ over the best performance between $M_{1}$ and $M_{2}$ (see Eq. 2). This metric assesses the weighted improvement gained by combining $M_{1}$ and $M_{2}$, compared to using only the best-performing model.
@@
We follow an iterative elimination approach to gain valuable insights into whether and to what extent combining multiple models enhances guessing effectiveness. We begin by combining passwords generated by all models and progressively remove those generated by the least effective model in terms of additional matched passwords. This process allows us to iteratively refine and identify effective combinations of models for any desired number of models to combine.

### v4

@@ Abstract
Recent advances in generative models have led to their application in password guessing, with the aim of replicating the complexity, structure, and patterns of human-created passwords. Despite their potential, inconsistencies and inadequate evaluation methodologies in prior research have hindered meaningful comparisons and a comprehensive, unbiased understanding of their capabilities. This paper introduces MAYA, a unified, customizable, plug-and-play benchmarking framework designed to facilitate the systematic characterization and benchmarking of deep generative password-guessing models in the context of trawling attacks. Using MAYA, we conduct a comprehensive assessment of six state-of-the-art DL-based models, which we re-implemented and adapted to ensure standardization, and two traditional ML-based approaches. Our evaluation spans eight real-world password datasets and covers an exhaustive set of advanced testing scenarios, totaling over 15,00015,\!000 compute hours. Our findings indicate that these models effectively capture different aspects of human password distribution and exhibit strong generalization capabilities. However, their effectiveness varies significantly with long and complex passwords. Through our evaluation, DL-based autoregressive models consistently outperform other deep learning approaches, demonstrating unique capabilities in generating accurate and complex guesses; meanwhile, ML-based approaches remain surprisingly highly competitive in many scenarios. Moreover, the diverse password distributions learned by the models enable a multi-model attack that outperforms the best individual model by an average of $\sim 7$ percentage points. By releasing MAYA, we aim to foster further research, providing the community with a new tool to consistently and reliably benchmark generative password-guessing models. Our framework is publicly available at https://github.com/williamcorrias/MAYA-Password-Benchmarking.git.
@@ 1 Introduction
Beyond these limitations, traditional Machine Learning password guessing techniques, such as Probabilistic Context-Free Grammars (PCFGs) [70, 38, 14, 72, 27] and Markov Models [46, 18, 25, 71], generate passwords based on statistical likelihood derived from real-world observed data. While these methods have demonstrated considerable effectiveness, they are fundamentally constrained by structural assumptions. Specifically, Markov Models, which rely on fixed-length n-grams, are inherently limited when modeling long-range dependencies and complex patterns. Likewise, PCFGs impose rigid template structures that constrain the diversity and flexibility of the generated password space.
@@
In recent years, research has increasingly focused on harnessing advancements in deep generative models to enhance password guessing techniques [33, 53, 50, 56, 48, 74, 69, 63]. Unlike traditional approaches, these models aim to learn and replicate the complexity, structure, and patterns of human-created passwords without relying on prior assumptions. Nevertheless, the current literature suffers from inconsistencies, inadequate evaluation practices, and a lack of rigorous, standardized characterization of models’ behaviors. Proposed approaches are often evaluated using inconsistent methodologies across studies and are tested in non-uniform, often simplistic settings, hindering meaningful comparisons and limiting a full understanding of their capabilities. As a result, critical aspects remain underexplored: how models behave across different settings, what types of passwords they tend to generate, how human-like their guesses are, whether they genuinely capture the underlying complexities of human password behavior, and where potential blind spots may lie. These limitations underscore a broader absence of systematic model characterization.
@@
To address these gaps, we introduce MAYA  [^1] The name ”MAYA” is inspired by A. Schopenhauer’s philosophical concept of ”the veil of Maya”, representing an illusion that obscures true reality. Similarly, our framework aims to unveil the potential of password-guessing generative models, which has remained obscured until now. , a unified and customizable plug-and-play benchmarking framework. While MAYA is specifically designed to support fair comparison and in-depth characterization and evaluation of DL-based generative password-guessing models in trawling attack scenarios, it builds upon a general, standardized, and rigorous evaluation methodology that can be applied across a wide range of guessing attacks, regardless of the underlying guessing approach. Using MAYA, we systematically evaluate eight generative models for password-guessing, comprising two ML-based and six DL-based approaches. Each deep generative model was re-implemented with standardized data preprocessing, dependencies, and configurations. We focus on trawling password attacks, as they represent the most general and widely applicable setting. Our experimental setup spans eight real-world password datasets for training and testing, and covers a wide range of evaluation scenarios, totaling over 15,00015,\!000 hours of computation.
@@
Our key findings include: (a) Increasing password length yields diminishing returns in reducing guessability; similarly, increasing the number of generated passwords offers reducing improvements in successful guesses, as models tend to exhaust easily guessable passwords early in the generation process. (b) Rule-based tools remain competitive only on smaller datasets, whereas traditional machine learning models perform consistently well, particularly on challenging ones. However, the best-performing deep learning model, on average, outperforms both. (c) Generative models do not always require large datasets to effectively model data distribution. While some approaches benefit from larger datasets, others exhibit minimal improvement or even a decrease in performance as the training data increases. (d) Models can successfully guess passwords even when trained on datasets from different communities and/or cultures compared to the test set, highlighting strong generalization capabilities and suggesting the existence of common structures underlying human-created passwords across disparate groups. (e) While all models demonstrate strong capabilities in guessing common and simple passwords and a decrease in performance for rarer passwords, autoregressive DL-based models and traditional ML-based approaches are the only ones that remain effective in guessing longer passwords and more complex patterns. (f) Different models learn generation functions with different probability distributions over the codomain, enabling their combination into a multi-model attack that outperforms individual models. (g) Generative models effectively capture various aspects of human-created passwords, generating high-quality and diverse passwords while minimizing mode failures. However, some models struggle to accurately replicate the length distribution.
@@
  • We develop MAYA, a fully customizable, plug-and-play framework for evaluating generative password-guessing models in trawling attack scenarios. To support reproducibility and encourage further research, we publicly release our code and data at https://github.com/williamcorrias/MAYA-Password-Benchmarking.git.
  • Leveraging MAYA, we characterize and benchmark eight state-of-the-art generative models across eight datasets and a diverse set of testing scenarios, addressing seven key research questions. Our evaluation spans over 15,00015,\!000 compute hours.
@@ 2.1 Lack of Consistency
Each existing approach adopts its own evaluation methodology, as models are neither trained nor assessed on the same data and under identical settings. These methodological inconsistencies encompass various factors, including data preprocessing algorithms, file encoding, vocabulary, maximum password length, the number of generated passwords, and the size of the training and testing datasets. For example, (1) PassGAN [33] compares models with different numbers of generated passwords; (2) VGPT2 [8] and FLA [48] define custom alphabets with ad-hoc special characters, whereas [56, 33, 53, 50] rely on the complete UTF-8 set; (3) PassFlow [50] is trained on a maximum of $300.000$ passwords and PassGPT [56] is trained only on unique passwords, while  [48, 33, 53, 8] adopt the standard 80-20 split without additional sampling; (4) FLA [48] filters passwords based on a minimum length, whereas [8, 56, 33, 53, 50] utilize a maximum length threshold (VGPT2 using 12, and the others 10). As a result, fair comparisons across studies are challenging. For instance, differing data preprocessing methods lead to distinct training and testing distributions, making direct comparison challenging. Likewise, evaluating models with different maximum password lengths introduces inherent bias, as shorter passwords are generally easier to guess. Analogous considerations apply to other settings as well. Such inconsistencies highlight the necessity for a standardized evaluation methodology to serve as a foundation for future research.
@@ 2.2 Lack of Rigorousness
Existing research often suffers from an insufficiently rigorous evaluation. Models are typically evaluated using overly simplistic metrics within restricted and similarly simplistic scenarios, such as the percentage of guessed passwords or the number of unique passwords generated  [33, 8, 50]. Moreover, part of the literature  [33, 50] relies only on widely used datasets like RockYou or LinkedIn. Consequently, the current literature offers an incomplete evaluation biased toward these simplistic settings, underscoring the need for more robust methodology incorporating diverse datasets, varied scenarios, and more complex settings that reflect a broader spectrum of real-world contexts.
@@ 2.4 Why Trawling Attacks
Password attacks are typically categorized based on the adversary’s scope: targeted and trawling. Targeted attacks aim to compromise specific user accounts by leveraging personal identifiable information (PII), and have gained increasing attention in recent years due to their growing effectiveness [42, 69, 51, 68, 30, 73]. In contrast, trawling attacks seek to guess as many user passwords as possible within a dataset, without targeting any specific individual. This attack model has long been the focus of password security research, with numerous techniques developed over time, including rule-based systems, PCFGs, Markov models, traditional neural networks, and deep generative models. In this work, we focus exclusively on trawling attacks, as they represent the most general and widely applicable class of password-guessing attacks, providing a broad evaluation framework. Targeted password-guessing attacks require different evaluation scenarios and datasets, and do not properly align with our RQs. Nevertheless, we plan to extend MAYA in future work to include them. Moreover, since targeted attacks can be viewed as trawling attacks conditioned on PII, our evaluation methodology can effectively support both lines of research.
@@ 3 MAYA
This section presents MAYA, our unified and plug-and-play benchmarking framework tailored for deep generative password-guessing models in the context of trawling attacks. The framework offers an intuitive environment for training and testing models with minimal setup, enabling the research community to move in a unified direction. It also includes a comprehensive set of experimental settings, providing thorough benchmarking and detailed characterization across multiple key metrics. MAYA features a highly modular and easily extendable architecture that enables the integration of new models and customized testing scenarios to support future research. While MAYA focuses on DL-based approaches, its underlying evaluation methodology (Section 3.1) provides a standardized testing environment that is applicable to all types of trawling password-guessing methods, ranging from traditional techniques to the latest approaches. MAYA currently implements six state-of-the-art deep generative password-guessing models and two traditional ML-based approaches (Section 3.2), eight real-world password datasets (Section 4), and a comprehensive set of testing scenarios aimed at answering seven key research questions (Section 3.3).
@@ 3.1.1 Standardized Data Preprocessing and Settings
We propose the following procedure to standardize data preprocessing: (1) open and read datasets using UTF-8 encoding while ignoring errors, (2) remove passwords that exceed the specified maximum length, contain non-ASCII characters, or include characters that are not in the provided vocabulary. (3) split the dataset following the standard $80\%$ training and $20\%$ testing ratios, (4) remove duplicates from the testing dataset, and (5) remove from the training dataset any overlap with the testing dataset. This approach ensures wide language compatibility through UTF-8 encoding, avoids double-counting by eliminating duplicates, and provides a fixed, consistent testing set across experiments. We note that existing works typically define the training set first, and then derive the testing set by removing any overlapping samples. This approach is undesirable, as varying the training set leads to changes in the testing set as well, thereby limiting the comparability across experiments. We further define our vocabulary to include all uppercase and lowercase letters, digits, and the following widely-accepted symbols: ˜!@#\$%ˆ&*(),.<>/?’"{}[$\|-$_=+;:
@@ 3.2 Models
TABLE I: Categorization of the selected approaches.
Model Trad. ML DL-Based Latent-Based Autoregressive FLA X X OMEN X X PassFlow X X PassGAN X X PassGPT X X PCFG X X X PLR-GAN X X VGPT2 X X
Table I shows the selected models and their respective categories used throughout the paper for reference. We selected eight generative models: two ML-based and six DL-based. Among them, four are latent-based approaches and three are autoregressive. We consider a model latent-based if it directly operates over a latent space $z$ [34].
Regarding DL-based models, we aimed to cover diverse architectural families: (1) FLA [48], an autoregressive approach based on an LSTM, (2) PassGAN [33], and (3) PLR-GAN [53], both GAN-based and thus latent-based, (4) PassFlow [50], a latent-based model built on a flow architecture, (5) VGPT2 [8], a latent-based approach combining a VAE with GPT2-derived encoder and decoder blocks, and (6) PassGPT [56], an autoregressive transformer model based on the GPT2 architecture. Our experiments also include two traditional ML-based password-guessing approaches: PCFG [70] relying on Probabilistic-Context-Free-Grammars, and OMEN [18] based on Markov Chains. This allows for a comprehensive comparative analysis between ML-based and DL-based methods, offering a complete overview and characterization of each approach.
Each DL-based model was fully or partially re-implemented to (1) standardize dependencies, and (2) enable a unified comparison framework. Specifically, we ported them to PyTorch 2.6.0 and adapted them to MAYA’s interface. These changes provide a common playground for fair comparisons and facilitate future research to work within the same testbed. We validated the accuracy of our implementations by comparing our results with those reported in the original papers, observing only negligible variations falling within the margin of error. Whenever possible, we also contacted the original authors for additional validation. We provide further implementation details in Appendix A.
@@
FLA [48] (Fast, Lean, and Accurate) was the first approach to apply neural networks to the password-guessing task. It is the only approach we selected based on a recurrent neural network, specifically an LSTM, allowing us to examine how autoregressive models perform compared to others, more recent generative architectures.
@@
PLR-GAN [53] is an enhanced version of PassGAN and represents the state-of-the-art for GAN-based methods. PLR-GAN adopts a Dynamic Password Guessing (DPG) strategy, which allows the model to adapt its guesses based on the distribution of successfully guessed passwords, increasing the likelihood of generating relevant guesses.
@@
PassGPT [56] proposes a GPT-2-based language model for password generation. Similarly to FLA, PassGPT employs an autoregressive generation process. However, GPT-2 is built upon an attention mechanism, which allows it to capture long-range dependencies more effectively.
PCFG
PCFG [70] learns a grammar -i.e., a set of rules describing password structures- from a dictionary of passwords. Given a structure (e.g., 4 letters followed by 4 digits), PCFG fills it using terminals. Thus, PCFG explicitly learns both the password patterns and the probabilities of each terminal within a pattern. The final probability of a password is calculated by multiplying the probability of the structure by the probabilities of the selected terminals. Passwords are generated in order of decreasing probability.
OMEN
OMEN [18] is a variant of the Markov n-gram model that generates passwords in descending probabilistic order, achieving higher performance than its traditional counterpart. Unlike PCFG, OMEN does not explicitly learn password patterns; instead, these are fixed and determined by the chosen n-gram window. What OMEN actually learns are the probabilities of character transitions within each window.
@@ 3.3 Research Questions
RQ2 - Are Deep Generative Models Truly Better than Traditional Tools?
Despite recent advancements in deep generative architectures, there is limited evidence indicating whether DL-based generative models outperform traditional tools and ML-based models, with only a few direct experimental comparisons conducted. We address this gap by conducting a comprehensive evaluation of generative models on 8 different datasets and comparing them to traditional tools such as John the Ripper (JtR) and Hashcat, offering a thorough, direct comparison between DL-based generative, ML-based generative, and rule-based methods.
@@
In real-world scenarios, attackers often have limited or partial access to leaks. Therefore, evaluating model performance across varying training data subset sizes is essential for gaining a clear understanding of the effectiveness of generative models. This RQ examines the ability of the models to reconstruct the full target data distribution starting from varying portions of the source dataset and execute a successful attack. We explore this capability by training models on up to seven different data subsets on four datasets.
@@
Simple passwords are easily guessed using traditional tools, whereas rare and complex passwords present a significantly greater challenge. The ability of generative models to arbitrarily sample from different areas of the password distribution offers the potential to guess even rare and complex passwords. We explore this RQ by dividing and categorizing the test dataset based on password frequency and assessing each model’s performance on the different subsets. Additionally, we analyze how models perform in guessing passwords of varying lengths and patterns.
@@
We explore the first part of this RQ using two metrics computed across all model pairs $(M_{1},M_{2})$: Jaccard Index and Mergeability Index. Jaccard Index measures the ratio between the intersection and union of the passwords generated by each model, providing a metric of diversity between $M_{1}$ and $M_{2}$ passwords. Additionally, we introduce the Mergeability Index, which quantifies the ratio between the marginal gain achieved by combining the outputs of $M_{1}$ and $M_{2}$ over the best performance between $M_{1}$ and $M_{2}$ (see Eq. 2). This metric assesses the weighted improvement gained by combining $M_{1}$ and $M_{2}$, compared to using only the best-performing model.
@@
We follow an iterative elimination approach to gain insights into whether and to what extent combining multiple models enhances guessing effectiveness. We begin by combining passwords generated by all models and progressively remove those generated by the least effective model in terms of additional matched passwords. This process allows us to iteratively refine and identify effective combinations of models for any desired number of models to combine.